OpenStudy (anonymous):

Simplify the following expression: x - 4y + 5z - 3x - y + 2z = ? 2x - 5y + 7z -2x - 5y - 7z -2x - 5y + 7z -2x -5y + 3z 2. . Evaluate the following expression using the values given: Find 5x - 3y - z if x = -2, y = 2, and z = -3 -19 19 13 -13

OpenStudy (anonymous):

-2x - 5y + 7z is the correct answer.

OpenStudy (tkhunny):

x - 4y + 5z - 3x - y + 2z = ? Go on a collection safari: x(1 - 3) + y(-4 - 1) + z(5+2)

OpenStudy (anonymous):

5*(-2) - 3*2 - (-3) = -10-6+3=-16+3=-13